40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 19 03:13:54 2011, kiwamu, Update, Electronics, REFL11 not working 1305799561ZgerKtrd.gif
    Reply  Fri May 20 11:37:44 2011, Koji, Update, Electronics, REFL11 fixed (REFL11 not working) REFL11_transimpedance.pdf
       Reply  Thu Jun 2 23:10:10 2011, Koji, Update, Electronics, REFL11 test results (Re: REFL11 fixed) REFL11_test.pdf
Message ID: 4756     Entry time: Fri May 20 11:37:44 2011     In reply to: 4747     Reply to this: 4782
Author: Koji 
Type: Update 
Category: Electronics 
Subject: REFL11 fixed (REFL11 not working) 

- Found the inductor which shunts the positive input of MAX4107 was not touching the ground.
This left the positive input level undetermined at DC. This was why MAX had been saturated.
The PCB has a cut, so it was surprising once the circuit worked.

- Resoldered the inductor to the ground. This made the circuit responding to the intensity-modulated beam.

- But the resonances and the notches were totally off, and the 200MHz oscillation has resurrected.

- Attached 40Ohm+22pF network between the neg-input of MAX and the gnd. This solved the oscillation.

- Made the tuning and the characterizations. The PD is on Kiwamu's desk and ready to go.

More to come later

Quote:
I took REFL11 out from the AS table for a health check because it wasn't working properly.

The symptoms were :

   - a big offset of ~ -3 V on the RF output. No RF signals.

   - The DC output seemed to be okay. It's been sensitive to light.

I did a quick check and confirmed that +/- 5V were correctly supplied to the op-amps.

It looks that the last stage (MAX4107) is saturated for some reasons. Need more inspections.

At the moment the REFL11 RFPD is on the bench of the Jenne laser.

 

Attachment 1: REFL11_transimpedance.pdf  40 kB  Uploaded Fri May 20 18:57:00 2011  | Hide | Hide all
REFL11_transimpedance.pdf
ELOG V3.1.3-